graph the equation, y = 8/7x + 3, using the slope and the y intercept
Answer by mananth(16946)

y=8/7x + 3
..
The y intercept point (0,3)
..
slope = rise/run
slope = 8/7 so from (0,3)move 8 units up the y axis and 7 units to the right. Mark the point(7,11).
.....
Similary move, from (0,3) 8 units down the y axis and 7 units to the left.(-5,7)
Mark these two points and draw the line.
...

...
m.ananth@hotmail.ca